Find the area of the shaded region. The graph to the right depicts IQ scores of​ adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15.+ + What color should I paint my x Homework and Tests - Laurel x Do Homework - Lauren John: X C Cet Homework Help With CHX ++ What color should I paint my X Homework and Tests - Laurel X Do Homework - Lauren John X C Cet Homework Help With CHX c mat+ + What color should I paint my x Homework and Tests - Laurel x Do Homework - Lauren John X C Cet Homework Help With CHX + c
Answer #1

solution

given : mean = 100 standard deviation = 15

z score = (X - mean )/ standard deviation : let X is a random variable

the area of the shaded region = P( 95 < X <125 ) =

=P( {95 - mean }/standard deviation < {X - mean }/standard deviation < {125 - mean }/standard deviation )

=P( -0.33 < z < 1.67)

= P(0< z< 1.67) + P(0 <z <0.33)

= 0.4525 + 0.1293 { from area of normal probabilty curve(table)}

=0.5818

so,the area of the shaded region= 0.5818